OpenStudy (anonymous):

The table and the equation shown each represent a different linear function. X Y 0 4 2 12 4 20 6 28 (y=7x+4) Write the difference in the rates of change for these two functions.

OpenStudy (mathstudent55):

The rate of change of a linear function is the slope. When you write a linear function's equation in the y = mx + b form, you can read the slope directly off the equation. It is m. What is the slope of the second function?

OpenStudy (anonymous):

I don't know. I'm behind in math right now and I haven't learned this stuff yet.

OpenStudy (welshfella):

Compare the 2 equations m = slope Y = mx + b Y l= 7x + 4 So for this equation m = ?

OpenStudy (anonymous):

I don't know. This stuff confuses me a lot. And like I said, I'm behind.

OpenStudy (welshfella):

What number in the second equation is in the same position as m in the first equation?

OpenStudy (anonymous):

7?

OpenStudy (welshfella):

Right - that indicates that 7 is the slope of the line (the graph of the equation will be a straight line.)

OpenStudy (anonymous):

Okay.

OpenStudy (welshfella):

The slope of the graph gives the rate of change of y compared with x. So if x increase ashes by 1 unit then y increases by 7units. if x increases by 2 then y will increase by 14 and so on.

OpenStudy (welshfella):

* ignore the 'ashes'. Lol

OpenStudy (anonymous):

Alright. :)

OpenStudy (welshfella):

Now the first function is written as a list of values of x and corresponding values of y You can work out the slope (rate of change) by looking at the relative differences in their values.

OpenStudy (anonymous):

Wait, is it 4?

OpenStudy (welshfella):

Note that as x goes from 0 to 2 y increases from 4 to 12 For every increase in 2 of x there's an increase in 8 of y So slope or rate of changes = 8 / 2 which = 4 Do you follow that OK?

OpenStudy (welshfella):

Yes. It's 4

OpenStudy (anonymous):

Yep. Thank you so much. :)

OpenStudy (welshfella):

So now u have the 2 rates of change Yw